Meeting Notes for 12 October

Present: Joel, Liz, Sudha, Mark, Ken, Marge (background)


  • The tutorials need updating, especially the graphs.
  • Several of the screencasts may need to be updated.
    • Ken is undecided about how to update the CMS screencasts. Some will depend on how the plot-saving function is finalized, requiring input from Tom. It's possible the 6 current screencasts can be replaced with one.
    • Sudha has an updated Performance screencast on i2u2-dev. Rolling it out to production is on her to-do list, but she asks for feedback first (link sent via email). Flux and Time-of-Flight still need attention.


  • The Shower Analysis step-by-step is on i2u2-dev. Sudha requests feedback before deploying to production.


  • Joel and Amber will work on identifying all of the missing data that Mihael discovered. After this data is replaced, we expect that we'll be able to successfully re-import the stream files and fix the plotting problems.
  • LIGO plot-saving still doesn't work
  • The "plot" button needs some UI attention

CIMA Improvements

No major updates.


Joel will upgrade the Hardware Enablement Stack (HWE) of the VMs this coming Friday evening. The upgrade tested well on i2u2-dev, so he'll do the remaining ones all at once while the site is down.

-- Main.JoelG - 2016-10-12


Topic revision: r1 - 2016-10-12, JoelG
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback